Expand e^2x in powers of (x-1), up to four terms
1
Expert's answer
2020-05-18T11:31:55-0400

"By\, Taylor's\, theorem,\, we\, know\, that \,
f(x) =""f(a)+(x-a)f'(a)+\frac{(x-a)^2}{2!}f''(a)+......"

"So, here f(x) =e^{2x}" and "a=1."

"e^{2x} =f(1)+(x-1)f'(1) +\frac{(x-1)^2}{2!}f''(1)+\frac{(x-1)^3}{3!}f'''(1)+......."

"f(1)=" "e^2 , f'(1)=2e^2, f''(1)=4e^2, f'''(1)=8e^2"

Putting the above values in the expansion of "e^{2x}" , we get

"e^{2x} = e^2+2(x-1)e^2+2(x-1)^2e^2+\frac{4}{3}(x-1)^3e^2" (expansion is done up to 4 terms)

Therefore, expansion of "e^{2x}" in powers of "(x-1)" up to 4 terms is

"e^{2x} = e^2+2(x-1)e^2+2(x-1)^2e^2+\frac{4}{3}(x-1)^3e^2 ."